Please answer the following:

Cost-Volume-Profit Relations: Missing Data
Following are data from 4 separate companies. Supply the missing data in each independent case

Case A Case B Case C Case D
Unit Sales 1,000 800 Answer Answer
Sales revenue $20,000 Answer Answer $60,000
Variable cost per unit $10 $2 $13 Answer
Contribution margin Answer $800 Answer Answer
Fixed Costs $7,500 Answer $80,000 Answer
Net income Answer $700 Answer Answer
Unit contribution margin Answer Answer Answer $19
Break-even point (units) Answer Answer 4,000 2,000
Margin of safety (units) Answer Answer 400 1,000
Answer #1
A B C D
Unit Sales 1000 800 4400 3000
Sales revenue $ 20,000.00 $ 2,400.00 $ 145,200.00 $ 60,000.00
Variable cost per unit $         10.00 $        2.00 $           13.00 $           1.00
Contribution margin $ 10,000.00 $    800.00 $   88,000.00 $ 57,000.00
Fixed Costs $   7,500.00 $    100.00 $   80,000.00 $ 38,000.00
Net income $   2,500.00 $    700.00 $     8,000.00 $ 19,000.00
Unit contribution margin $         10.00 $        1.00 $           20.00 $        19.00
Break-even point (units) 750 100 4000 2,000
Margin of safety (units) 250 700 400 1000

2 с 10 800 =(C3*C5)+C6 =D10+D11 =(D3*D5)+D6 13 3 Unit Sales 1000 4 Sales revenue 20000 5 Variable cost per unit 10 6 Contribu

General Formulas
Break Even Point = Fixed Cost / Contribution Margin p.u
Margin of safety = Actual Sales - Break Even Point
Cont. Margin = Sales - Variable Costs
Cont. Margin p.u = Contribution Margin / Unit sales